OpenStudy (lxelle):

Show 2ae^1/2a - 4e^1/2a = 2 as x=2+e^-1/2x

OpenStudy (thomas5267):

Could you use brackets to clarify the expression?

OpenStudy (lxelle):

\[2ae ^{a/2} - 4e^{a/2} =2\] as \[x = 2+ e^{-x/2}\]

OpenStudy (phi):

can you post a copy of the original question?

OpenStudy (lxelle):

I was doing it halfway, then i couldnt show that x = ______ equation.

OpenStudy (phi):

ok. \[ 2a e^\frac{a}{2}- 4e^\frac{a}{2}=2 \\ 2e^\frac{a}{2}(a-2) = 2 \\ e^\frac{a}{2}(a-2) =1 \\ a-2 = e^{-\frac{a}{2}} \]

OpenStudy (phi):

and finally a= 2 + exp(-a/2)

OpenStudy (lxelle):

thankss. could you help me with ques 5ii?

OpenStudy (phi):

did you find "a" ?

OpenStudy (lxelle):

oh yes a =3.

OpenStudy (lxelle):

so (2x+1) ( 2x^2-3x+3) <0

OpenStudy (phi):

I assume you mean \[ (2x+1) ( 2x^2-3x+3) <0 \] for that expression to be negative we need either 1. 2x+1 < 0 and (2x^2-3x+3) is positive or 2. vice versa

OpenStudy (lxelle):

Okay.

OpenStudy (lxelle):

< 0 means expression has to be negative, is it?

OpenStudy (phi):

yes, stuff < 0 means stuff is less than 0 (i.e. negative) the roots of (2x^2-3x+3) are complex that means this parabola does not cross the x-axis the 2 in front of x^2 makes it concave up (smile shape) and as a check, the vertex is at -(-3)/(2*2) = 3/4. that means it's min y value is at x=3/4 and it is 2*9/16 - 9/4 + 3 = 15/8 so the parabola is always positive that means we need the other factor to always be negative because we need minus time plus to get a minus

OpenStudy (lxelle):

I seee. so we take 2x+1 <0? is it?

OpenStudy (phi):

yes

OpenStudy (lxelle):

nicee thanks!!another ques. no 4!!!

OpenStudy (lxelle):

just want you to check my answer.

OpenStudy (lxelle):

can you see?

OpenStudy (lxelle):

.

OpenStudy (phi):

as a check to Q5, here is a plot, showing the curve goes negative for x< -0.5

OpenStudy (lxelle):

oh cool! okayy!! thanks!

OpenStudy (lxelle):

there was once you showed me a similar working to that ques no. 4. I was quite confused about this. how is sin2x/ (cos2x +1) = sinx /cosx ? D:

OpenStudy (phi):

yes, it looks ok until you "jump" to the answer use: sin(2x) = 2 sinx cosx cos(2x) = cos^2 x - sin^2(x) which can be rewritten using cos^2 + sin^2 = 1 i.e. cos^2 x = 1 - sin^2 x, so cos(2x) = (1-sin^2 x) - sin^2 x = 1 - 2 sin^2x

OpenStudy (phi):

now you can work on \[ \frac{\sin 2x}{1 - \cos 2x } \]

OpenStudy (phi):

you should get cos x / sin x = cot x

OpenStudy (lxelle):

I'm lost! :( So, dy/dx = sin 2x / (1-cos2x) dy/dx = 2 sinxcosx / (1 - sin^2 x) is it?

OpenStudy (phi):

cos2x = 1 - 2 sin^2x so you second line should read dy/dx = 2 sinxcosx / (1 -( 1 - 2 sin^2 x))

OpenStudy (phi):

\[ \frac{dy}{dx} = \frac{2 \sin x \cos x}{1 - 1 + 2 \sin^2 x} \]

OpenStudy (lxelle):

what should i do next? i kept going back to the previous pattern.

OpenStudy (phi):

1-1 = 0 \[ \frac{dy}{dx} = \frac{2 \sin x \cos x}{1 - 1 + 2 \sin^2 x} \\ = \frac{2 \sin x \cos x}{2 \sin^2 x} \]

OpenStudy (phi):

the denominator starts as \[ 1 - \cos 2x \] we have the identity \[ \cos 2x = 1 - 2\sin^2 x \] so the bottom becomes \[ 1 - ( 1 - 2\sin^2 x) \\ = 1 - 1 + 2\sin^2 x \\ = 2 \sin^2 x \]

OpenStudy (lxelle):

okay what's next?

OpenStudy (phi):

simplify \[ \frac{dy}{dx}= \frac{2 \sin x \cos x}{2 \sin^2 x} \]

OpenStudy (lxelle):

cannot do. D: I kept getting back sin2x/ 1- cox2x

OpenStudy (phi):

if the problem was \[ \frac{dy}{dy}= \frac{2 w \ z}{2 w^2} \] can you simplify that ?

OpenStudy (lxelle):

cos x / sinx?

OpenStudy (phi):

yes. \[ \frac{2 \sin x \cos x}{2 \sin x \sin x} = \frac{\cos x }{\sin x } \]

OpenStudy (lxelle):

OHH. I see it now! thankss!

OpenStudy (lxelle):

Could Pls check my solution for me. The question is (1/ 1-x) + (2 / 1+2x) - (4 / 2+x)